# Learning to Mask and Permute Visual Tokens for Vision Transformer Pre-Training

### v1

@@ Abstract
The use of self-supervised pre-training has emerged as a promising approach to enhance the performance of visual tasks such as image classification. In this context, recent approaches have employed the Masked Image Modeling paradigm, which pre-trains a backbone by reconstructing visual tokens associated with randomly masked image patches. This masking approach, however, introduces noise into the input data during pre-training, leading to discrepancies that can impair performance during the fine-tuning phase. Furthermore, input masking neglects the dependencies between corrupted patches, increasing the inconsistencies observed in downstream fine-tuning tasks. To overcome these issues, we propose a new self-supervised pre-training approach, named Masked and Permuted Vision Transformer (MaPeT), that employs autoregressive and permuted predictions to capture intra-patch dependencies. In addition, MaPeT employs auxiliary positional information to reduce the disparity between the pre-training and fine-tuning phases. In our experiments, we employ a fair setting to ensure reliable and meaningful comparisons and conduct investigations on multiple visual tokenizers, including our proposed $k$-CLIP which directly employs discretized CLIP features. Our results demonstrate that MaPeT achieves competitive performance on ImageNet, compared to baselines and competitors under the same model setting. Source code and trained models are publicly available at: https://github.com/aimagelab/MaPeT.
@@ 1 Introduction
Self-supervised pre-training models have achieved remarkable success in boosting the performance of Transformer-based architectures in Computer Vision. Inspired by the BERT model devlin2019bert , the Masked Image Modeling (MIM) pre-training objective bao2022beit has become widely adopted in literature as a powerful self-supervision method for vision tasks. In particular, this pre-training objective involves masking random image patches and reconstructing the corrupted visual input. While several recent studies have refined the MIM approach chen2022context ; el2021large ; he2022masked , there has been little exploration of alternative pre-training objectives in the visual domain. In contrast, in the field of Natural Language Processing (NLP), several methods song2020mpnet ; yang2019xlnet have surmounted the BERT pre-training objective with the introduction of different paradigms that aim to solve the drawbacks of previous methods.
@@
Drawing inspiration from NLP, we investigate a permutation-based pre-training strategy, which we term Permuted Image Modeling (PIM). This approach autoregressively predicts permuted image patches maintaining contextual bi-directionality without corrupting any part of the input. Despite offering an improvement over the standard MIM-based objective, the autoregressive technique of PIM reduces the amount of positional information available for each prediction. To tackle this issue, we propose a Masked and Permuted pre-training solution for Vision Transformers (MaPeT) which leverages auxiliary position information as input during pre-training, thereby allowing the model to access the positional information of every image patch.
@@
In addition to the pre-training objective, a crucial aspect of self-supervised vision pre-training is the design of visual targets, used as supervisory signal. While some works have employed low-level and hand-crafted visual features huang2022contrastive ; liu2022swin ; wei2022masked , the dominant approach is to employ discrete visual tokens to reconstruct the corrupted visual input bao2022beit ; chen2022context ; el2021large ; peng2022beit . In this context, although BEiT bao2022beit initially employed DALL-E ramesh2021zero visual tokens, its performance has been surpassed by VQ-KD, proposed in BEiT v2 peng2022beit . In particular, VQ-KD employs an encoder-decoder architecture that reconstructs CLIP semantic features and is directly trained on ImageNet-1k deng2009imagenet , requiring retraining to achieve satisfactory results on other datasets. In contrast to previous works, we propose $k$-CLIP, a novel discrete tokenizer for generating visual tokens that can directly employ CLIP features without requiring training an ad hoc discrete autoencoder.
@@ 2 Related Work
Self-supervised learning. Several solutions have been introduced in the last years to effectively pre-train vision-based architectures via self-supervised learning, initially based on different pretext tasks doersch2015unsupervised ; gidaris2018unsupervised ; noroozi2016unsupervised ; wang2015unsupervised ; zhang2016colorful and then exploiting contrastive learning paradigms chen2020simple ; chen2021exploring ; grill2020bootstrap ; he2020momentum ; oord2018representation . The advent of Vision Transformer models (ViT) dosovitskiy2021image has pushed towards the introduction of new increasingly sophisticated self-supervised pre-training strategies bao2022beit ; caron2021emerging ; chen2020generative ; chen2021empirical . In this context, motivated by the great success in NLP, some attempts have been made to effectively adapt the Masked Language Modeling paradigm devlin2019bert ; liu2019roberta and its auto-regressive variant brown2020language in the Computer Vision domain, either by directly predicting pixels chen2020generative , image patches dosovitskiy2021image , or discrete visual tokens bao2022beit . In particular, the recently proposed BEiT approach bao2022beit effectively performs pre-training via image patch masking and predicts the discretized labels ramesh2021zero of masked patches. Since the introduction of BEiT, many subsequent methods based on similar pre-training strategies have been presented chen2022context ; dong2023peco ; el2021large ; he2022masked ; wei2022masked ; wei2022mvp ; xie2022simmim . Some of these methods chen2022context ; el2021large ; he2022masked have introduced an encoder-decoder architecture to separate uncorrupted encoded information from masked tokens, which are employed directly as input to the decoder. Inspired by findings from NLP literature song2020mpnet ; yang2019xlnet , our proposed method aims to overcome the limitations of Masked Image Modeling in self-supervised pre-training.
@@
Visual targets. In the domain of self-supervised learning, visual targets are a set of specific objectives that are used as the supervisory signal to pre-train visual models. Based on the type of signal, the targets can be categorized into low-level visual features, hand-crafted features, and visual tokens. Some recent studies fang2022corrupted ; he2022masked ; huang2022contrastive ; liu2022swin ; xie2022simmim have utilized pixel information as the low-level supervisory signal for self-supervised pre-training. In contrast, wei2022masked have used HOG hand-crafted features to reconstruct the masked visual input. More recently, visual targets based on CLIP radford2021learning have been utilized with remarkable success, either by directly employing CLIP features MILAN2022 ; wei2022mvp ; zhang2022cae or training a discrete tokenizer peng2022beit to reconstruct the semantic features encoded by CLIP. In our method, we propose a discrete visual tokenizer based on CLIP features which offers a novel approach to visual pre-training without the need for specific training over a particular dataset.
@@ 3 Preliminaries
Figure 1: (a) Masked Image Modeling (MIM). (b) Permuted Image Modeling (PIM). (c) Masked and Permuted pre-training for Vision Transformers (MaPeT). While MIM reconstructs visual tokens from randomly masked image patches, PIM autoregressively predicts tokens associated with permuted image patches. MaPeT uses PIM to capture intra-patch dependency and takes auxiliary position information as input to ensure that the model sees a full sequence of patches at each target position.
@@
Image patches. In this study, we adopt ViT dosovitskiy2021image as the backbone network for our architecture. ViT splits an image into a sequence of 2D image patches, which are linearly projected to the model embedding space and elaborated through multiple attention blocks. Given an input image $\bm{x}\in\mathbb{R}^{H\times W\times C}$, this is mapped into a sequence of $N$ square patches $\{\bm{x}_{i}^{p}\}_{i=1}^{N}$, where $\bm{x}_{i}^{p}\in\mathbb{R}^{P\times P\times C}$ is the $i$-th patch of the input image. Subsequently, a linear layer is applied to each flattened patch to project it to the input dimensionality of the model $D$, outputting the patch embeddings $\{\bm{x}_{i}\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{D}$, which are then added to the learnable 1D positional embeddings $\bm{E}_{\text{pos}}\in\mathbb{R}^{N\times D}$.
The ViT dosovitskiy2021image encoder consists of $L$ identical layers of Transformer blocks, where the output embeddings of the last layer represent the encoded representations of the $N$ input image patches. In our experiments, we consider an image dimension of $224\times 224$ with a patch dimension $P$ of $16\times 16$, constituting an input sequence of $14\times 14=196$ patch embeddings.
@@
Masked Image Modeling (MIM). Inspired by the Masked Language Modeling strategy utilized in BERT devlin2019bert , the MIM paradigm is a pre-training technique for vision tasks that aims to recover visual information from a corrupted input image. This is achieved by randomly masking a portion of the image patches and predicting the visual tokens related to the corrupted region of the input, as depicted in Fig. 1 (a). Like Masked Language Modeling, MIM has some inherent disadvantages. Firstly, a mask token $M$ is introduced during pre-training and never used during fine-tuning, leading to a discrepancy in the pre-training and fine-tuning phases. Secondly, given the masked tokens $\bm{\bar{x}}$ and the uncorrupted context $\bm{\tilde{x}}$, the probability $p(\bm{\bar{x}}\mid\bm{\tilde{x}})$ is typically factorized, assuming the independence of reconstructed patches. To address these issues, the NLP literature has investigated a permutation-based variant yang2019xlnet that can reduce the disadvantages of standard Masked Language Modeling. These results motivate us to investigate the application of this strategy to vision tasks.
@@
While MIM preserves full positional information of each image patch, during the prediction process, PIM can access the contextual and positional information only of the $t-1$ patches that precede $z_{t}$, as shown in Eq. 1. Given this lack of full positional information, PIM introduces an input discrepancy between pre-training and fine-tuning, underscoring the necessity for a pre-training technique that combines the advantages of both MIM and PIM while mitigating their respective limitations.
@@ 4 Proposed Method
4.1 Masked and Permuted Pre-Training for Vision Transformers
Drawing inspiration from NLP literature song2020mpnet ; yang2019xlnet , we propose a novel pre-training methodology called Masked and Permuted Vision Transformer (MaPeT) which builds on the strengths of both MIM and PIM to enhance performance on vision tasks. In particular, our approach overcomes the independence assumption of reconstructed patches, thus capturing intra-patch dependencies more effectively. Moreover, MaPeT incorporates auxiliary position embeddings during pre-training, enabling the model to access position information for all patches, thereby resolving the pre-training fine-tuning discrepancy introduced by PIM. Fig. 1 (c) shows an overview of the MaPeT approach.
Given a permutation $\bm{z}$ and a cutting point $c$, MaPeT can predict the visual token associated with an input patch embedding $\bm{x}_{z_{t}}$ by leveraging the content and position of the preceding $\bm{x}_{z_{<t}}$, as well as the position of the subsequent target embeddings $\bm{x}_{z_{>t}}$. To this end, we introduce the concept of learnable masked token $M\in\mathbb{R}^{D}$, which is used to express the positional information of $\bm{x}_{z_{>c}}$. By repeating $N-c$ times the token $M$, we obtain $\{M_{i}\}_{i=c+1}^{N}$ identical masked tokens, which are then summed to the positional embedding $\{E_{\text{pos}}^{i}\}_{i=z_{c+1}}^{z_{N}}$ of each target $\bm{x}_{z_{>c}}$. These resulting position-aware masked tokens $\bm{M}_{\text{pos}}=\{M_{c+1}+E_{\text{pos}}^{z_{c+1}},\dots,M_{N}+E_{\text{pos}}^{z_{N}}\}$ are concatenated to the input sequence of patch embeddings $\bm{H}^{0}$ thus obtaining the augmented input $\bm{H}^{0}_{M}=[\bm{H}^{0},\bm{M}_{\text{pos}}]$. Note that the positional embeddings are permuted according to the same permutation $\bm{z}$ applied to the patches. For convenience, we introduce ${\bm{M}_{z_{\geq t}}}$ which represents the subset $\{M_{\text{pos}}^{i}\}_{i=\text{max}(1,\;t-c)}^{N-c}$. Intuitively, when $t\leq c$, we have that ${\bm{M}_{z_{\geq t}}}$ comprises all $\bm{M}_{\text{pos}}$. Conversely, when $t>c$, ${\bm{M}_{z_{\geq t}}}$ only includes the position-aware masked tokens related to $\bm{x}_{z_{\geq t}}$.
@@ 4.1 Masked and Permuted Pre-Training for Vision Transformers
Figure 2: Overview of our MaPeT pre-training (a), with content stream attention (b), which follows the standard self-attention mechanism, and query stream attention (c), which lacks information about the content of the patch embedding $\bm{x}_{z_{t}}$ whose visual token $v_{z_{t}}$ is to be predicted. The blue and green masks in (a) are the content and query attention masks employed in the two-stream self-attention.
@@
Two-stream self-attention pre-training. Since the target patch embeddings $\bm{x}_{z>c}$ follow the permuted order, the next predicted patch can occur in any position. As a consequence, masking the attention matrix of a ViT encoder, instead of corrupting the input like in MIM, makes the backbone architecture non-trivial. To implement PIM on a ViT backbone, we adopt the two-stream self-attention mechanism introduced by XLNet yang2019xlnet . Specifically, our model consists of two attention streams: a query stream and a content stream. The query stream $g_{\theta}(\bm{x}_{z_{<t}},{\bm{M}_{z_{\geq t}}})$ accesses the content of the previous patches in the permuted sequence at a given position $t$. However, it does not access the content of $\bm{x}_{z_{t}}$, only viewing the position ${\bm{M}_{z_{\geq t}}}$ of the subsequent patches. In contrast, the content stream $h_{\theta}(\bm{x}_{z_{\leq t}},{\bm{M}_{z_{>t}}})$ encodes the content of both the previous elements in the sequence and the element in position $z_{t}$, with positional information ${\bm{M}_{z_{>t}}}$ of the remaining patches. The input of the first query stream layer $\bm{g}^{(0)}$ consists of the masked elements $\bm{M}_{\text{pos}}$, which encode the position of the target patch embeddings $\bm{x}_{z_{>c}}$. Instead, the input of the first content stream layer $\bm{h}^{(0)}$ is the augmented input sequence $\bm{H}^{0}_{M}$. Formally, for each Transformer layer $l$ with $l=1,\ldots,L$, the attention mechanism of both content and query streams can be defined as follows:
$h^{(l)}_{z_{t}}\xleftarrow{}\mathsf{Attention}(\bm{Q}=h^{(l-1)}_{z_{t}},\bm{KV}=(\mathbf{h}^{(l-1)}_{\mathbf{{z_{\leq t}}}},\mathbf{h}^{(l-1)}_{\mathbf{{M_{z_{>t}}}}});\,\theta),$ (3)
$g^{(l)}_{z_{t}}\xleftarrow{}\mathsf{Attention}(\bm{Q}=g^{(l-1)}_{z_{t}},\bm{KV}=(\mathbf{h}^{(l-1)}_{\mathbf{{z_{<t}}}},\mathbf{h}^{(l-1)}_{\mathbf{{M_{z_{\geq t}}}}});\,\theta),$
@@
where $\bm{Q}$, $\bm{K}$, and $\bm{V}$ are respectively queries, keys, and values of the attention operator. Both query and content streams use separate attention masks to limit the visible contextual and positional information for each patch. During pre-training, the output of the query stream is used as the model output. At the fine-tuning step, the query stream is dropped, and only the content stream is used, returning to the standard ViT backbone. Fig. 2 shows an illustration of the content stream (b), the query stream (c), and how they are integrated into our MaPeT architecture.
@@
Attention masking. In order to limit the number of visible patches in the content and query attention operations, as described in Eq. 3, MaPeT leverages attention masking. In particular, the permutation $z$ influences the creation of two distinct attention masks: one for the content attention stream and one for the query attention stream. The content mask guarantees that only patch embeddings $\bm{x}_{z_{\leq t}}$ and positional tokens $\bm{M}_{z_{>t}}$ are visible to the patch embedding $\bm{x}_{z_{t}}$ in the content stream. On the other hand, the query mask ensures that only patch embeddings $\bm{x}_{z_{<t}}$ and positional tokens $\bm{M}_{z_{\geq t}}$ are visible to $\bm{x}_{z_{t}}$ in the query stream.

### v2

@@ Abstract
The use of self-supervised pre-training has emerged as a promising approach to enhance the performance of many different visual tasks. In this context, recent approaches have employed the Masked Image Modeling paradigm, which pre-trains a backbone by reconstructing visual tokens associated with randomly masked image patches. This masking approach, however, introduces noise into the input data during pre-training, leading to discrepancies that can impair performance during the fine-tuning phase. Furthermore, input masking neglects the dependencies between corrupted patches, increasing the inconsistencies observed in downstream fine-tuning tasks. To overcome these issues, we propose a new self-supervised pre-training approach, named Masked and Permuted Vision Transformer (MaPeT), that employs autoregressive and permuted predictions to capture intra-patch dependencies. In addition, MaPeT employs auxiliary positional information to reduce the disparity between the pre-training and fine-tuning phases. In our experiments, we employ a fair setting to ensure reliable and meaningful comparisons and conduct investigations on multiple visual tokenizers, including our proposed $k$-CLIP which directly employs discretized CLIP features. Our results demonstrate that MaPeT achieves competitive performance on ImageNet, compared to baselines and competitors under the same model setting. We release an implementation of our code and models at https://github.com/aimagelab/MaPeT.
journal: Computer Vision and Image Understanding
@@ 1 Introduction
**footnotetext: Both authors contributed equally to this research.
@@
Self-supervised pre-training models have achieved remarkable success in boosting the performance of Transformer-based architectures in Computer Vision. Inspired by the BERT model (Devlin et al., 2019), the Masked Image Modeling (MIM) pre-training objective (Bao et al., 2022) has become widely adopted in literature as a powerful self-supervision method for vision tasks. In particular, this pre-training objective involves masking random image patches and reconstructing the corrupted visual input. While several recent studies have refined the MIM approach (Chen et al., 2023; He et al., 2022), there has been little exploration of alternative pre-training objectives in the visual domain. In contrast, in the field of Natural Language Processing (NLP), several methods (Song et al., 2020; Yang et al., 2019) have surpassed the BERT pre-training objective with the introduction of different paradigms that aim to solve the drawbacks of previous methods.
@@
Drawing inspiration from NLP, we investigate a permutation-based pre-training strategy, which we term Permuted Image Modeling (PIM). This approach autoregressively predicts permuted image patches maintaining contextual bi-directionality without corrupting any part of the input. Despite offering an improvement over the standard MIM-based objective, the autoregressive PIM technique reduces the amount of positional information available for each prediction. To tackle this issue, we propose a Masked and Permuted pre-training solution for Vision Transformers (MaPeT) which leverages auxiliary position information as input during pre-training, thus allowing the model to access the positional information of each image patch.
@@
In addition to the pre-training objective, a crucial aspect of self-supervised vision pre-training is the design of visual targets, used as supervisory signal. While some works have employed low-level and hand-crafted visual features (Huang et al., 2023; Liu et al., 2022; Wei et al., 2022a), the dominant approach is to employ discrete visual tokens to reconstruct the corrupted input (Bao et al., 2022; Chen et al., 2023; El-Nouby et al., 2021; Peng et al., 2022). In this context, although BEiT (Bao et al., 2022) initially employed DALL-E (Ramesh et al., 2021) visual tokens, its performance has been surpassed by VQ-KD, proposed in BEiT v2 (Peng et al., 2022). In particular, VQ-KD employs an encoder-decoder architecture that reconstructs CLIP features and is directly trained on ImageNet-1k (Deng et al., 2009), requiring retraining to achieve satisfactory results on other datasets. In contrast to previous works, we propose $k$-CLIP, a novel discrete tokenizer for generating visual tokens that can directly employ CLIP features without requiring training an ad hoc discrete autoencoder.
Fig. 1: (a) Masked Image Modeling (MIM). (b) Permuted Image Modeling (PIM). (c) Masked and Permuted pre-training for Vision Transformers (MaPeT). While MIM reconstructs visual tokens from randomly masked image patches, PIM autoregressively predicts tokens associated with permuted image patches. MaPeT uses PIM to capture intra-patch dependency and takes auxiliary position information as input to ensure that the model sees a full sequence of patches at each target position.
@@ 2 Related Work
Self-supervised learning. Several solutions have been introduced in the last years to effectively pre-train vision-based architectures via self-supervised learning, initially based on different pretext tasks (Doersch et al., 2015; Gidaris et al., 2018; Noroozi and Favaro, 2016; Wang and Gupta, 2015; Zhang et al., 2016) and then exploiting contrastive learning paradigms (Chen et al., 2020b; Chen and He, 2021; Grill et al., 2020; He et al., 2020; Oord et al., 2018). The advent of Vision Transformer models (ViT) (Dosovitskiy et al., 2021) has pushed towards the introduction of new increasingly sophisticated self-supervised pre-training strategies (Bao et al., 2022; Caron et al., 2021; Chen et al., 2020a, 2021). In this context, motivated by the great success in NLP, some attempts have been made to effectively adapt the Masked Language Modeling paradigm (Devlin et al., 2019; Liu et al., 2019) and its auto-regressive variant (Brown et al., 2020) in the Computer Vision domain, either by directly predicting pixels (Chen et al., 2020a), image patches (Dosovitskiy et al., 2021), or discrete visual tokens (Bao et al., 2022). In particular, the recently proposed BEiT approach (Bao et al., 2022) effectively performs pre-training via image patch masking and predicts the discretized labels (Ramesh et al., 2021) of masked patches. Since the introduction of BEiT, many subsequent methods based on similar pre-training strategies have been presented (Chen et al., 2023; Dong et al., 2023; El-Nouby et al., 2021; He et al., 2022; Tian et al., 2022, 2023; Wei et al., 2022a, b; Xie et al., 2022; Zhang et al., 2023). Some of these methods (Chen et al., 2023; El-Nouby et al., 2021; He et al., 2022) have introduced an encoder-decoder architecture to separate uncorrupted encoded information from masked tokens, which are employed directly as input to the decoder. Zhang et al. (2023) introduced HiViT, a simplified hierarchical Vision Transformer that combines the advantages of ViT and Swin (Liu et al., 2022) models while being more efficient for Masked Image Modeling. (Tian et al., 2023) proposed integrally pre-trained Transformer pyramid networks, which jointly optimize a HiViT-based network backbone and a feature pyramid neck to minimize the transfer gap between pre-training and downstream tasks. Additionally, (Tian et al., 2022) investigated alternative learning objectives for token-based pre-training, providing insights into effective design principles beyond traditional masking approaches. Inspired by findings from NLP literature (Song et al., 2020; Yang et al., 2019), our proposed method aims to overcome the limitations of Masked Image Modeling in self-supervised pre-training.
@@
Visual targets. In the domain of self-supervised learning, visual targets are a set of specific objectives that are used as the supervisory signal to pre-train visual models. Based on the type of signal, the targets can be categorized into low-level visual features, hand-crafted features, and visual tokens. Some recent studies (Fang et al., 2023; He et al., 2022; Huang et al., 2023; Liu et al., 2022; Xie et al., 2022) have utilized pixel information as the low-level supervisory signal for self-supervised pre-training. In contrast, (Wei et al., 2022a) have used HOG hand-crafted features to reconstruct the masked visual input. Recently, visual targets based on CLIP (Radford et al., 2021) have been used with remarkable success, either by directly employing CLIP features (Hou et al., 2022; Wei et al., 2022b; Zhang et al., 2022) or training a discrete tokenizer (Peng et al., 2022) to reconstruct the semantic features encoded by CLIP. In our method, we propose a discrete visual tokenizer based on CLIP features which offers a novel approach to visual pre-training without the need for specific training over a particular dataset.
@@ 3 Preliminaries
Image patches. In this study, we adopt ViT (Dosovitskiy et al., 2021) as the backbone network for our architecture. ViT splits an image into a sequence of 2D image patches, which are linearly projected to the model embedding space and elaborated through multiple attention blocks. Given an input image $\bm{x}\in\mathbb{R}^{H\times W\times C}$, this is mapped into a sequence of $N$ square patches $\{\bm{x}_{i}^{p}\}_{i=1}^{N}$, where $\bm{x}_{i}^{p}\in\mathbb{R}^{P\times P\times C}$ is the $i$-th patch of the input image. Subsequently, a linear layer is applied to each flattened patch to project it to the input dimensionality of the model $D$, outputting the patch embeddings $\{\bm{x}_{i}\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{D}$, which are then added to the learnable 1D positional embeddings $\bm{E}_{\text{pos}}\in\mathbb{R}^{N\times D}$.
The ViT (Dosovitskiy et al., 2021) encoder consists of $L$ identical layers of Transformer blocks, where the output embeddings of the last layer represent the encoded representations of the $N$ input image patches. In our experiments, we consider an image dimension of $224\times 224$ with a patch dimension $P$ of $16\times 16$, constituting an input sequence of $14\times 14=196$ patch embeddings.
@@
Masked Image Modeling (MIM). Inspired by the Masked Language Modeling strategy utilized in BERT (Devlin et al., 2019), the MIM paradigm is a pre-training technique for vision tasks that aims to recover visual information from a corrupted input image. This is achieved by randomly masking a portion of the image patches and predicting the visual tokens related to the corrupted region of the input, as depicted in Fig. 1 (a). Like Masked Language Modeling, MIM has some inherent disadvantages. Firstly, a mask token $M$ is introduced during pre-training and never used during fine-tuning, leading to a discrepancy in the pre-training and fine-tuning phases. Secondly, given the masked tokens $\bm{\bar{x}}$ and the uncorrupted context $\bm{\tilde{x}}$, the probability $p(\bm{\bar{x}}\mid\bm{\tilde{x}})$ is typically factorized, assuming the independence of reconstructed patches. To address these issues, the NLP literature has investigated a permutation-based variant (Yang et al., 2019) that can reduce the disadvantages of standard Masked Language Modeling. These results motivate us to investigate the application of this strategy to vision tasks.
@@
While MIM preserves full positional information of each image patch, during the prediction process, PIM can access the contextual and positional information only of the $t-1$ patches preceding $z_{t}$, as shown in Eq. 1. Given this lack of full positional information, PIM introduces an input discrepancy between pre-training and fine-tuning, underscoring the need for a pre-training technique that combines the advantages of both MIM and PIM while mitigating their respective limitations.
@@ 4 Proposed Method
Fig. 2: Overview of our MaPeT pre-training (a), with content stream attention (b), which follows the standard self-attention mechanism, and query stream attention (c), which lacks information about the content of the patch embedding $\bm{x}_{z_{t}}$ whose visual token $v_{z_{t}}$ is to be predicted. The blue and green masks in (a) are the content and query attention masks employed in the two-stream self-attention.
@@
4.1 Masked and Permuted Pre-Training
Drawing inspiration from NLP literature (Song et al., 2020; Yang et al., 2019), we propose a novel pre-training methodology called Masked and Permuted Vision Transformer (MaPeT) which builds on the strengths of both MIM and PIM to enhance performance on vision tasks. In particular, our approach overcomes the independence assumption of reconstructed patches, thus capturing intra-patch dependencies more effectively. Moreover, MaPeT incorporates auxiliary position embeddings during pre-training, enabling the model to access position information for all patches, thereby resolving the pre-training fine-tuning discrepancy introduced by PIM. Fig. 1 (c) shows an overview of the MaPeT approach.
Given a permutation $\bm{z}$ and a cutting point $c$, MaPeT can predict the visual token associated with an input patch embedding $\bm{x}_{z_{t}}$ by leveraging the content and position of the preceding $\bm{x}_{z_{<t}}$, as well as the position of the subsequent target embeddings $\bm{x}_{z_{>t}}$. To this end, we introduce the concept of learnable masked token $M\in\mathbb{R}^{D}$, which is used to express the positional information of $\bm{x}_{z_{>c}}$. By repeating $N-c$ times the token $M$, we obtain $\{M_{i}\}_{i=c+1}^{N}$ identical masked tokens, which are then summed to the positional embedding $\{E_{\text{pos}}^{i}\}_{i=z_{c+1}}^{z_{N}}$ of each target $\bm{x}_{z_{>c}}$. These resulting position-aware masked tokens $\bm{M}_{\text{pos}}=\{M_{c+1}+E_{\text{pos}}^{z_{c+1}},\dots,M_{N}+E_{\text{pos}}^{z_{N}}\}$ are concatenated to the input patch embeddings $\bm{H}^{0}$ thus obtaining the augmented input $\bm{H}^{0}_{M}=[\bm{H}^{0},\bm{M}_{\text{pos}}]$. Note that the positional embeddings are permuted according to the same permutation $\bm{z}$ applied to the patches. For convenience, we introduce ${\bm{M}_{z_{\geq t}}}$ which represents the subset $\{M_{\text{pos}}^{i}\}_{i=\text{max}(1,\;t-c)}^{N-c}$. Intuitively, when $t\leq c$, we have that ${\bm{M}_{z_{\geq t}}}$ comprises all $\bm{M}_{\text{pos}}$. Conversely, when $t>c$, ${\bm{M}_{z_{\geq t}}}$ only includes the position-aware masked tokens related to $\bm{x}_{z_{\geq t}}$.
@@ 4.1 Masked and Permuted Pre-Training
Two-stream self-attention pre-training. Since the target patch embeddings $\bm{x}_{z>c}$ follow the permuted order, the next predicted patch can occur in any position. As a consequence, masking the attention matrix of a ViT encoder, instead of corrupting the input like in MIM, makes the backbone architecture non-trivial. To implement PIM on a ViT backbone, we adopt the two-stream self-attention mechanism introduced by XLNet (Yang et al., 2019). Specifically, our model consists of two attention streams: a query stream and a content stream. The query stream $g_{\theta}(\bm{x}_{z_{<t}},{\bm{M}_{z_{\geq t}}})$ accesses the content of the previous patches in the permuted sequence at a given position $t$. However, it does not access the content of $\bm{x}_{z_{t}}$, only viewing the position ${\bm{M}_{z_{\geq t}}}$ of the subsequent patches. In contrast, the content stream $h_{\theta}(\bm{x}_{z_{\leq t}},{\bm{M}_{z_{>t}}})$ encodes the content of both the previous elements in the sequence and the element in position $z_{t}$, with positional information ${\bm{M}_{z_{>t}}}$ of the remaining patches. The input of the first query stream layer $\bm{g}^{(0)}$ consists of the masked elements $\bm{M}_{\text{pos}}$, which encode the position of the target patch embeddings $\bm{x}_{z_{>c}}$. Instead, the input of the first content stream layer $\bm{h}^{(0)}$ is the augmented input sequence $\bm{H}^{0}_{M}$. Formally, for each Transformer layer $l$ with $l=1,\ldots,L$, the attention mechanism of both content and query streams can be defined as:
$h^{(l)}_{z_{t}}\xleftarrow{}\mathsf{Attention}(\bm{Q}=h^{(l-1)}_{z_{t}},\bm{KV}=(\bm{h}^{(l-1)}_{{z_{\leq t}}},\bm{h}^{(l-1)}_{{\bm{M}_{z_{>t}}}});\,\theta),$ (3)
$g^{(l)}_{z_{t}}\xleftarrow{}\mathsf{Attention}(\bm{Q}=g^{(l-1)}_{z_{t}},\bm{KV}=(\bm{h}^{(l-1)}_{{z_{<t}}},\bm{h}^{(l-1)}_{{\bm{M}_{z_{\geq t}}}});\,\theta),$
@@
where $\bm{Q}$, $\bm{K}$, and $\bm{V}$ are respectively queries, keys, and values of the attention operator. Both query and content streams use separate attention masks to limit the visible contextual and positional information for each patch. During pre-training, the output of the query stream is used as the model output. At the fine-tuning step, the query stream is dropped, and only the content stream is used, returning to the standard ViT backbone. Fig. 2 shows the content stream (b), the query stream (c), and how they are integrated into our MaPeT architecture.
@@
Attention masking. To limit the number of visible patches in the content and query attention operations, as described in Eq. 3, MaPeT leverages attention masking. In particular, the permutation $z$ influences the creation of two distinct attention masks: one for the content attention stream and one for the query attention stream. The content mask guarantees that only patch embeddings $\bm{x}_{z_{\leq t}}$ and positional tokens $\bm{M}_{z_{>t}}$ are visible to the patch embedding $\bm{x}_{z_{t}}$ in the content stream. On the other hand, the query mask ensures that only patch embeddings $\bm{x}_{z_{<t}}$ and positional tokens $\bm{M}_{z_{\geq t}}$ are visible to $\bm{x}_{z_{t}}$ in the query stream.